# PlantBert: An Open Source Language Model for Plant Science

### v1

@@ Abstract
The rapid advancement of transformer-based language models has catalyzed breakthroughs in biomedical and clinical natural language processing; however, plant science remains markedly underserved by such domain-adapted tools. In this work, we present PlantBert, a high-performance, open-source language model specifically tailored for extracting structured knowledge from plant stress-response literature. Built upon the DeBERTa architecture—known for its disentangled attention and robust contextual encoding—PlantBert is fine-tuned on a meticulously curated corpus of expert-annotated abstracts, with a primary focus on lentil (Lens culinaris) responses to diverse abiotic and biotic stressors. Our methodology combines transformer-based modeling with rule-enhanced linguistic post-processing and ontology-grounded entity normalization, enabling PlantBert to capture biologically meaningful relationships with precision and semantic fidelity. The underlying corpus is annotated using a hierarchical schema aligned with the Crop Ontology, encompassing molecular, physiological, biochemical, and agronomic dimensions of plant adaptation. PlantBert exhibits strong generalization capabilities across entity types and demonstrates the feasibility of robust domain adaptation in low-resource scientific fields.By providing a scalable and reproducible framework for high-resolution entity recognition, PlantBert bridges a critical gap in agricultural NLP and paves the way for intelligent, data-driven systems in plant genomics, phenomics, and agronomic knowledge discovery. Our model is publicly released to promote transparency and accelerate cross-disciplinary innovation in computational plant science.
@@ 1 Introduction
To address this gap, we introduce PlantBert, an open-source, domain-adapted language model built upon the DeBERTa architecture [6]. PlantBert is fine-tuned on a meticulously annotated corpus of lentil (Lens culinaris) abstracts, focusing on biotic and abiotic stress conditions. It leverages a hybrid pipeline that combines transformer modeling, part-of-speech (POS) augmentation, rule-based post-processing, and ontology-aligned term normalization. PlantBert is designed to facilitate accurate entity recognition and support structured knowledge extraction for plant biology and agricultural research [5, 10].
@@
  • We present PlantBert, the first DeBERTa-based language model specifically fine-tuned for plant stress-response literature.
@@
  • We publicly release the PlantBert model and dataset, enabling reproducibility and providing a foundation for future agricultural NLP research.
In the following sections, we contextualize our approach within related literature, outline our data and model design pipeline, and demonstrate PlantBert’s performance across key NER benchmarks within the plant science domain.
@@ 3 Methodology
Figure 1: Overview of the PlantBert pipeline. (a) Agricultural and plant literature is collected and filtered for domain relevance. (b) Experts manually annotate plant stress-related entities using a structured schema. (c) The annotated data is used to fine-tune a DeBERTa-based model, followed by post-processing for syntactic adjustments and heuristic corrections. (d) The final model performs high-resolution entity labeling to support downstream applications.
@@ 3.1 Transformer-Based Model Selection
We adopted a model-driven approach to identify the most suitable transformer backbone for domain-specific named entity recognition (NER) in plant science. Our selection process was guided by evaluating multiple pre-trained architectures commonly used in the biomedical and scientific NLP communities, including BERT, DistilBERT, BioBERT, and DeBERTa. Based on prior benchmarks and architectural strengths, DeBERTa was ultimately selected as the foundation for PlantBert due to its disentangled attention mechanism and improved positional encoding, which are particularly effective for complex, structured text domains such as plant science literature.
@@ 3.3 Evaluation of Annotation Performance
3.4 PlantBert Architecture and Training Pipeline
PlantBert is a domain-adapted transformer model built upon the DeBERTa architecture, fine-tuned for token-level classification. The implementation pipeline included the following components:
@@ 3.5 Post-Processing and Ontology Alignment
This hybrid pipeline, combining statistical modeling with domain knowledge, forms the core of PlantBert’s high-precision NER capabilities.
@@ 4 Results and Discussion
We conduct a comparative evaluation of PlantBert against widely used transformer-based models for Named Entity Recognition (NER), including BERT [5], DistilBERT [14], BioBERT [10], and DeBERTa [6]. All models were fine-tuned and evaluated on the same domain-specific corpus using identical training configurations to ensure a fair and unbiased comparison.
As shown in Table 1, PlantBert achieves the highest scores across all macro and weighted F1 metrics, including a macro F1-score of 0.8269 and a weighted F1-score of 0.9549, outperforming all baseline models by substantial margins. Although DeBERTa reports the highest overall accuracy (0.8905), PlantBert’s edge on macro-averaged metrics indicates a more balanced performance across all entity types—particularly important in domain-specific corpora where entity distributions are often highly skewed.
The improvement in both macro precision (0.8278) and macro recall (0.8366) demonstrates PlantBert’s ability to make accurate predictions even for underrepresented entities. These results are crucial for real case scenarios in plant science, where many meaningful terms—such as gene variants, ecological traits, or cultivar-specific expressions—appear infrequently in the literature. Traditional models tend to bias predictions toward frequent classes, but PlantBert’s fine-grained recall capacity makes it a more robust solution in high-stakes scenarios where information loss is unacceptable.
Moreover, the weighted F1-score of 0.9549 confirms that PlantBert does not sacrifice performance on dominant entity classes in favor of rare ones. Instead, it achieves a superior balance, suggesting an advanced understanding of domain-specific language, even when the surface forms of entities are ambiguous or semantically subtle.
@@
Beyond performance metrics, these results highlight an important insight: PlantBert’s improvements stem not from architectural novelty but from careful domain-aware pretraining and corpus selection. Its development involved curating a comprehensive plant science corpus that includes journal articles, databases, and taxonomic resources, ensuring broad coverage of terminologies used in plant breeding, genomics, phenomics, and agronomy. This specialization directly translates to more meaningful and actionable NER outputs in applied plant science pipelines.
@@
PlantBert 0.8838 0.8278 0.9610 0.8366 0.9522 0.8269 0.9549
In summary, PlantBert sets a new benchmark for NER in plant science by achieving high recall and precision across both frequent and rare entities. Its performance supports the broader argument for domain-specific language models, especially in scientific domains where vocabulary, structure, and semantics deviate significantly from general corpora.
@@ 5 Conclusion
In this study, we introduced PlantBert, a domain-adapted transformer model engineered for named entity recognition in plant science. Built upon DeBERTa and enriched through expert annotation, ontology alignment, and post-processing heuristics, PlantBert consistently outperforms general and biomedical baselines across all key NER metrics. Designed initially for Lens culinaris stress literature, PlantBert’s architecture is inherently scalable—supporting future extensions to additional crops, stressors, and languages. Its robust performance underscores the importance of full-stack domain adaptation, from corpus curation to symbolic refinement. To support reproducibility and interdisciplinary research, PlantBert and its annotated dataset are publicly available at: https://huggingface.co/PHENOMA/PlantBert. We anticipate this resource will serve as a shared foundation for plant scientists, computational linguists, and agronomic researchers working toward data-driven discovery in crop science. Future improvements to PlantBert may include enhanced knowledge graph construction by leveraging NER outputs to populate structured representations of plant science concepts, as well as multilingual adaptation to broaden its utility across diverse linguistic contexts. Incorporating continual learning from new literature and evolving domain ontologies could further maintain its relevance and accuracy over time.

### v3

# PlantDeBERTa: An Open Source Language Model for Plant Science

@@ Abstract
The rapid advancement of transformer-based language models has catalyzed breakthroughs in biomedical and clinical natural language processing; however, plant science remains markedly underserved by such domain-adapted tools. In this work, we present PlantDeBERTa, a high-performance, open-source language model specifically tailored for extracting structured knowledge from plant stress-response literature. Built upon the DeBERTa architecture—known for its disentangled attention and robust contextual encoding—PlantDeBERTa is fine-tuned on a meticulously curated corpus of expert-annotated abstracts, with a primary focus on lentil (Lens culinaris) responses to diverse abiotic and biotic stressors. Our methodology combines transformer-based modeling with rule-enhanced linguistic post-processing and ontology-grounded entity normalization, enabling PlantDeBERTa to capture biologically meaningful relationships with precision and semantic fidelity. The underlying corpus is annotated using a hierarchical schema aligned with the Crop Ontology, encompassing molecular, physiological, biochemical, and agronomic dimensions of plant adaptation. PlantDeBERTa exhibits strong generalization capabilities across entity types and demonstrates the feasibility of robust domain adaptation in low-resource scientific fields.By providing a scalable and reproducible framework for high-resolution entity recognition, PlantDeBERTa bridges a critical gap in agricultural NLP and paves the way for intelligent, data-driven systems in plant genomics, phenomics, and agronomic knowledge discovery. Our model is publicly released to promote transparency and accelerate cross-disciplinary innovation in computational plant science.
@@ 1 Introduction
To address this gap, we introduce PlantDeBERTa, an open-source, domain-adapted language model built upon the DeBERTa architecture [6]. PlantDeBERTa is fine-tuned on a meticulously annotated corpus of lentil (Lens culinaris) abstracts, focusing on biotic and abiotic stress conditions. It leverages a hybrid pipeline that combines transformer modeling, part-of-speech (POS) augmentation, rule-based post-processing, and ontology-aligned term normalization. PlantDeBERTa is designed to facilitate accurate entity recognition and support structured knowledge extraction for plant biology and agricultural research [5, 10].
@@
  • We present PlantDeBERTa, the first DeBERTa-based language model specifically fine-tuned for plant stress-response literature.
@@
  • We publicly release the PlantDeBERTa model and dataset, enabling reproducibility and providing a foundation for future agricultural NLP research.
In the following sections, we contextualize our approach within related literature, outline our data and model design pipeline, and demonstrate PlantDeBERTa’s performance across key NER benchmarks within the plant science domain.
@@ 3 Methodology
Figure 1: Overview of the PlantDeBERTa pipeline. (a) Agricultural and plant literature is collected and filtered for domain relevance. (b) Experts manually annotate plant stress-related entities using a structured schema. (c) The annotated data is used to fine-tune a DeBERTa-based model, followed by post-processing for syntactic adjustments and heuristic corrections. (d) The final model performs high-resolution entity labeling to support downstream applications.
@@ 3.1 Transformer-Based Model Selection
We adopted a model-driven approach to identify the most suitable transformer backbone for domain-specific named entity recognition (NER) in plant science. Our selection process was guided by evaluating multiple pre-trained architectures commonly used in the biomedical and scientific NLP communities, including BERT, DistilBERT, BioBERT, and DeBERTa. Based on prior benchmarks and architectural strengths, DeBERTa was ultimately selected as the foundation for PlantDeBERTa due to its disentangled attention mechanism and improved positional encoding, which are particularly effective for complex, structured text domains such as plant science literature.
@@ 3.3 Evaluation of Annotation Performance
3.4 PlantDeBERTa Architecture and Training Pipeline
PlantDeBERTa is a domain-adapted transformer model built upon the DeBERTa architecture, fine-tuned for token-level classification. The implementation pipeline included the following components:
@@ 3.5 Post-Processing and Ontology Alignment
This hybrid pipeline, combining statistical modeling with domain knowledge, forms the core of PlantDeBERTa’s high-precision NER capabilities.
@@ 4 Results and Discussion
We conduct a comparative evaluation of PlantDeBERTa against widely used transformer-based models for Named Entity Recognition (NER), including BERT [5], DistilBERT [14], BioBERT [10], and DeBERTa [6]. All models were fine-tuned and evaluated on the same domain-specific corpus using identical training configurations to ensure a fair and unbiased comparison.
As shown in Table 1, PlantDeBERTa achieves the highest scores across all macro and weighted F1 metrics, including a macro F1-score of 0.8269 and a weighted F1-score of 0.9549, outperforming all baseline models by substantial margins. Although DeBERTa reports the highest overall accuracy (0.8905), PlantDeBERTa’s edge on macro-averaged metrics indicates a more balanced performance across all entity types—particularly important in domain-specific corpora where entity distributions are often highly skewed.
The improvement in both macro precision (0.8278) and macro recall (0.8366) demonstrates PlantDeBERTa’s ability to make accurate predictions even for underrepresented entities. These results are crucial for real case scenarios in plant science, where many meaningful terms—such as gene variants, ecological traits, or cultivar-specific expressions—appear infrequently in the literature. Traditional models tend to bias predictions toward frequent classes, but PlantDeBERTa’s fine-grained recall capacity makes it a more robust solution in high-stakes scenarios where information loss is unacceptable.
Moreover, the weighted F1-score of 0.9549 confirms that PlantDeBERTa does not sacrifice performance on dominant entity classes in favor of rare ones. Instead, it achieves a superior balance, suggesting an advanced understanding of domain-specific language, even when the surface forms of entities are ambiguous or semantically subtle.
@@
Beyond performance metrics, these results highlight an important insight: PlantDeBERTa’s improvements stem not from architectural novelty but from careful domain-aware pretraining and corpus selection. Its development involved curating a comprehensive plant science corpus that includes journal articles, databases, and taxonomic resources, ensuring broad coverage of terminologies used in plant breeding, genomics, phenomics, and agronomy. This specialization directly translates to more meaningful and actionable NER outputs in applied plant science pipelines.
@@
PlantDeBERTa 0.8838 0.8278 0.9610 0.8366 0.9522 0.8269 0.9549
In summary, PlantDeBERTa sets a new benchmark for NER in plant science by achieving high recall and precision across both frequent and rare entities. Its performance supports the broader argument for domain-specific language models, especially in scientific domains where vocabulary, structure, and semantics deviate significantly from general corpora.
@@ 5 Conclusion
In this study, we introduced PlantDeBERTa, a domain-adapted transformer model engineered for named entity recognition in plant science. Built upon DeBERTa and enriched through expert annotation, ontology alignment, and post-processing heuristics, PlantDeBERTa consistently outperforms general and biomedical baselines across all key NER metrics. Designed initially for Lens culinaris stress literature, PlantDeBERTa’s architecture is inherently scalable—supporting future extensions to additional crops, stressors, and languages. Its robust performance underscores the importance of full-stack domain adaptation, from corpus curation to symbolic refinement. To support reproducibility and interdisciplinary research, PlantDeBERTa and its annotated dataset are publicly available at: https://huggingface.co/PHENOMA/PlantDeBERTa. We anticipate this resource will serve as a shared foundation for plant scientists, computational linguists, and agronomic researchers working toward data-driven discovery in crop science. Future improvements to PlantDeBERTa may include enhanced knowledge graph construction by leveraging NER outputs to populate structured representations of plant science concepts, as well as multilingual adaptation to broaden its utility across diverse linguistic contexts. Incorporating continual learning from new literature and evolving domain ontologies could further maintain its relevance and accuracy over time.